40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 27 16:47:53 2015, ericq, Update, IOO, MC2 -> MCL Actuator TF mc2_2_mcl.pngMC2_2_MCL_TF.txt.zip
    Reply  Thu Jul 30 02:06:20 2015, Ignacio, Update, IOO, MC2 <-> MCL Actuator TF fitted TF_BODE.pngMC2_2_MCL.pngTF_INV_BODE.pngMCL_2_MC2.png
    Reply  Wed Aug 12 22:33:36 2015, Ignacio, Update, IOO, Re-measured MC2 -> MCL TF bode_TF.pngCoherence.png
    Reply  Fri Aug 28 20:15:49 2015, rana, Update, IOO, MC2 -> MCL Actuator TF 
       Reply  Mon Aug 31 07:49:11 2015, ericq, Update, IOO, MC2 -> MCL Actuator TF 
          Reply  Mon Aug 31 14:15:23 2015, Ignacio, Update, IOO, Measured the MC_F whitening poles/zeroes Data.ziptf.png
Message ID: 11550     Entry time: Mon Aug 31 14:15:23 2015     In reply to: 11548
Author: Ignacio 
Type: Update 
Category: IOO 
Subject: Measured the MC_F whitening poles/zeroes 

I measured the 15 Hz zero and the 150 Hz pole for the whitening filter channels of the Generic Pentek board in the IOO rack. The table below gives these zero/pole pairs for each of the 8 channels of the board.

channel zero [Hz] pole [Hz] Chan
1 15.02 151.05 C1:ASC-POP_QPD_YAW
2 15.09 150.29 C1:ASC-POP_QPD_PIT
3 14.98 150.69 C1:ASC-POP_QPD_SUM
4 14.91 147.65 C1:ALS-TRX
5 15.03 151.19 C1:ALS-TRY
6 15.01 150.51 ---
7 14.95 150.50 C1:IOO-MC_L
8 15.03 150.93 C1:IOO-MC_F

Here is a plot of one of the measured transfer functions,

and the measured data is attached here: Data.zip

EQ: I've added the current channels going through this board. 

More importantly, I found that the jumpers on channel one (QPD X) were set to no whitening, in contrast to all other channels. Thus, the POP QPD YAW signals we've been using for who knows how long have been distorted by dewhitening. This has now been fixed. 

Hence, the current state of this board is that the first whitening stage is disabled for all channels and the second stage is engaged, with the above parameters. 

Attachment 1: Data.zip  245 kB  Uploaded Mon Aug 31 16:07:00 2015
ELOG V3.1.3-